1. What is the effect of each of the following factors (increase, decrease, or has no effect) on the value of the molar mass of the liquid:

a. If the thermometer used to measure the temperature of water reads 1.5 °C higher than the true temperature.

b. If the atmospheric pressure is mistakenly recorded less than the actual value.

c. The liquid in the flask was not vaporized completely,

d. If the student used 2 mL of the liquid more than in the procedure.

2. A volatile liquid of mass 1.286 g occupies 538 mL at 98.7 °C and 748 mmHg. Calculate the number of moles and the molar mass of the liquid.


1
Expert's answer
2021-01-12T07:02:53-0500

1) a. Increase

b. Increase

c. Increase

d. Decrease

2) PV = n RT

P = 748mmHg × "\frac{1atm}{760mmHg}" = 0.98 atm

n = "\frac{mass}{molar.mass}" = "\frac{m}{M}"

PV = "\frac{m}{M}" RT

P = 0.98atm , m = 1.286g , V = 0.538L R = 0.08206 L-atm/K-mol , T = 273+ 98.7 = 371.7K

M = "\frac{mRT}{PV}"

M = "\frac{1.286×0.08206×371.7}{0.98×0.538}"

M = 74.39 g/mol

no. of moles ( n ) = m/M

n = 1.286/74.39

n = 0.017 mol